nitially, the switch in the circuit below is closed. If the circuit is still closed what is the current in the resistance? in A, 2 significant figures

lake iu the current to decrease by factor equal e=2.71828 (E significant figures +5v 3 400 uF 202 200
Answer #1

Here if switch is closed for a long time then no current passes through capacitor then entire current passes through the resistor.

Hence current, i =V/R = 5/2 = 2.5 A .